Suppose approximately 56% of all students study after midnight.
For random sample of n=3 students, let X be the number of students that study after midnight.
Complete the following table for the binomial distribution of X.
Keep at least 4 decimal places.
X P(X)
0
1
2
3
Determine the mean and standard deviation of X.
Keep at least 4 decimal places.
The expected value of X=

The standard deviation of X=

Answer by Boreal(15235)

X P(X)
0 .44^3=0.0852
1 3C1*.56^1*.44^2=0.3252
2 3C2*0.56^2*0.44=0.4140
3 0.56^3=0.1756
They add to 1.0, as they must.
E(x)=np=3*0.56=1.68
V(x)=np(1-p)=1.68*0.44=0.7392
sd= sqrt (V)=0.8598